Operations Report – May 5th

Crew 235 Operations Report 5-MAY-2021

SOL: 3

Name of person filing report: Atila Meszaros

Non-nominal systems: 1 EVA suit, Science Dome door, and inner airlock door of the RAM

Notes on non-nominal systems: Shannon used a screwdriver, a drill, a socket set, and a sledgehammer to fix both doors!

Generator: Working nominally.

Hours run: 0 hours

From what time last night: the generator didn’t run

To what time this morning: the generator didn’t run

List any additional daytime hours when the generator was run: N/A

Solar— SOC 98% at 18:23 (Before generator is run at night)

Notes on power system: Nothing to report

Diesel Reading – 47%

Station Propane Reading – 80%

Outpost Propane Reading – 1. 71% 2. 80%

Water (loft tank): 50 gallons

Water Meter: 150994.0 units

Water (static tank): 350 gallons

Static to Loft Pump used – yes

Water in GreenHab: 0 gallons

Water in ScienceDome: 0 gallons

Toilet tank emptied: No

Sojourner rover used: No, in town

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Spirit rover used: No, in town

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Opportunity rover used: No, in town

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Curiosity rover used: No, in town

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Perseverance rover used: Yes

Hours: 209.2

Beginning charge: 100%

Ending charge: 97%

Currently charging: Yes

Notes on rovers: None

ATV’s Used: Yes

Reason for use: EVA#3 and EVA#4

Oil Added? No

ATV Fuel Used: N/A

# Hours the ATVs were used today: 0

Notes on ATVs: ATVs will be filled with ethanol-free gas to be ready for tomorrow’s EVA.

HabCar used and why, where?: Nothing to report

CrewCar used and why, where?: Nothing to report

General notes and comments: Nothing to report

Summary of internet: Working nominally.

Summary of suits and radios: All radios nominal, all functional suits currently charging.

Summary of Hab operations: Mouse found in the morning and released during EVA #3.

Summary of GreenHab operations: Door not quite fixed.

Summary of ScienceDome operations: Door fixed.

Summary of RAM operations: Door fixed.

Summary of any observatory issues: The rotation problem was a battery issue.

Summary of health and safety issues: Nothing to report

Questions, concerns, and requests to Mission Support: Thank you for being with us tonight!